Concern:

Have the diesel trucks proven to be cost saving?

Information:

Diesel fuel is 12 cents per gallon cheaper than gasoline

Average Truck gets 10 miles per gallon

Diesel trucks cost more to repair and maintain than gasoline trucks.

Data has been collected on 20 trucks with the following information: age, miles driven and type.


“Tony,” said Mike as he put the data in Table 1 in front of Sanchez, “this data shows that, on average the 6 diesel trucks cost $21,156 in R&M expense whereas the 14 gasoline trucks averaged only $13,269 in R&M expense. I am also surprised that simple correlations show a negative relationship between miles driven and age and between miles driven and type of truck.” Tony looked up at Mike and replied, “Mike can I borrow this data. I have a young college graduate that I just hired and she is very bright. Let me ask her to look at this data and see what she comes up with.”


Required:
Assume that you are an analyst working for Tony Sanchez. He has asked you to analyze this data and to address the questions below. Prepare a business memo that addresses his issues.

  1. On average, how much extra does it cost to repair a diesel truck compared to a gasoline Truck?
Truck # Expense Miles Years Type
1      21,100    100,625 1 1
2      23,548    108,750 2 1
3      18,560      83,438 2 1
4      19,525      77,500 3 1
5      22,435      88,125 3 1
6      21,768      85,938 3 1
7      10,345      95,625 1 0
8      11,575      91,875 1 0
9      14,240      98,125 2 0
10      13,300      88,438 2 0
11      14,105      90,625 2 0
12      13,050      81,875 2 0
13      17,150      95,625 3 0
14      11,230      76,849 2 0
15      10,490      92,145 1 0
16      16,645      86,940 3 0
17      16,219      91,257 3 0
18      13,565      85,432 2 0
19         9,245      84,688 1 0
20      14,600      80,625 3 0
Answer #1

We can calculate expense per mile on both the trucks . We can clearly see that first 6 trucks are diesel trucks .

Total distance travelled by those 6 trucks = 544376

Total expense incurred on those 6 trucks = $119236

Average cost per mile = 119236/544736 = $0.2189 per mile

Total distance travelled by 14 gasoline trucks = 1240124 miles

Total expense incurred on those 14 gasoline trucks = $185759

Expenses per mile = 185759/1240124 = $0.1498 per mile

Extra cost incurred for repair of diesel truck in comparison to gasoline truck = 0.2189-0.1498 = $0.0691.

Total extra cost incurred = 0.0691x544376 = $37616.